Prescott Football Manufacturing had the following operating results for 2019: sales = $30,724; cost of goods sold = $21,938; depreciation expense = $3,589; interest expense = $599; dividends paid = $893. At the beginning of the year, net fixed assets were $20,365, current assets were $2,061, and current liabilities were $4,997. At the end of the year, net fixed assets were $23,242, current assets were $4,568, and current liabilities were $3,253. The tax rate for 2019 was 22 percent.

a.       What is net income for 2019? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

b.       What is the operating cash flow for 2019? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

c.       What is the cash flow from assets for 2019? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

d.       Assume no new debt was issued during the year. What is the cash flow to creditors for 2019? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

e.       Assume no new debt was issued during the year. What is the cash flow to stockholders for 2019? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
Answer #1

a.       What is net income for 2019? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

Parameter Linkage $
Sales A          30,724
[-] COGS B          21,938
[-] Depreciation C            3,589
EBIT D = A - B - C            5,197
[-] Interest E               599
EBT F = D - E            4,598
[-] Taxes G = 22% x F            1,012
Net income H = F - G            3,586

b.       What is the operating cash flow for 2019? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

Operating Cash Flow (OCF) = EBIT + Depreciation − Tax = 5,197 + 3,589 - 1,012 = $ 7,774

c.       What is the cash flow from assets for 2019? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

Cash Flow from Asset = OCF – Net Capital Spending (NCS) – Changes in Net Working Capital (NWC)

OCF: 7,774 from part (b) above

NCS = Changes in Net Fixed Asset (NFA) + Depreciation = (Ending NFA−Beginning NFA)+Depreciation = (23,242 − 20,365) + 3,589 =  6,466

NWC = Current Assets – Current Liabilities

Change in NWC = Ending NWC−Beginning NWC = (4,568 - 3,253) - (2,061 - 4,997) = 4,251

Hence, cash flow from assets = 7,774 - 6,466 - 4,251 = - 2,943

d.       Assume no new debt was issued during the year. What is the cash flow to creditors for 2019? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

the cash flow to creditors for 2019 = Interest paid - new debt issued = 599 - 0 = 599

e.       Assume no new debt was issued during the year. What is the cash flow to stockholders for 2019? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

Cash flow to stockholders = cash flow from assets - cash flow to creditor = - 2,943 - 599 = - 3,542